3. A necessary and sufficient condition for a matrix (square) A to be invertible is that
a.A is not equal zero
b.|A|≠0
c.|A|>0
d.A<0

4. Given that x + 2y + 3z =1, 3x + 2y + z = 4, x + 3y + 2z = 0. What is x,y and z?
a.(7/4,-3/4,1/4)
b.(5/4,-2/4,1/4)
c.(1,-3/4,1/4)
d.(1,-2,3)
1
Expert's answer
2019-04-09T10:39:20-0400

3b - it is a known theorem

4a

"x=\frac{\begin{vmatrix}
 1 & 2 & 3 \\
 4 & 2 & 1\\
 0&3&2
\end{vmatrix}}{\begin{vmatrix}
 1 & 2 & 3 \\
 3 & 2 & 1\\
1&3&2
\end{vmatrix}}=\frac{4+36-16-3}{4+2+27-6-12-3}=\frac{7}{4}"







"y=\frac{\begin{vmatrix}
 1 & 1 & 3 \\
 3 & 4 & 1\\
 1&0&2
\end{vmatrix}}{\begin{vmatrix}
 1 & 2 & 3 \\
 3 & 2 & 1\\
1&3&2
\end{vmatrix}}=\frac{8+1-12-6}{4+2+27-6-12-3}=-\frac{3}{4}"




"z=\frac{\begin{vmatrix}
 1 & 2 & 1 \\
 3 & 2 & 4\\
 1&3&0
\end{vmatrix}}{\begin{vmatrix}
 1 & 2 & 3 \\
 3 & 2 & 1\\
1&3&2
\end{vmatrix}}=\frac{8+9-2-12}{4+2+27-6-12-3}=\frac{1}{4}"